Consider the function f(x)=3−6x2,  −4≤x≤2f(x)=3-6x2,  -4≤x≤2.

The absolute maximum value is:
and this occurs at: x=

The absolute minimum value is:
and this occurs at: x=
Answer #1

foc) = 3-6x² -4EX L2 3. f(x) = 3-6x? [-4,2) sia) = -122 critical point falls EO) -12x50=0 ало of x) = -12 So f(0) = -1270 h